# Program evaluation with remotely sensed outcomes

### v2

@@ 2.2 Main assumption: Stability
We formalize this causal setting via three assumptions. The first is standard.
Assumption 2.1 (Experimental unconfoundedness).
@@
    SUTVA: $Y=DY(1)+(1-D)Y(0)$ almost surely.
@@
In many empirical applications involving RSVs, such as Examples 2.1 and 2.2, Assumption 2.1 is satisfied by design: experimental units are chosen as aggregates without spillovers, e.g. villages, and the treatment is randomly assigned for these experimental units.
Under Assumption 2.1, if we were to observe the outcome in the experimental sample, the ATE could be identified using standard arguments. However, the outcome is not observed in the experiment; instead, we have an RSV.
We resolve this measurement issue by leveraging the observational sample. Intuitively, Assumption 2.2 allows us to learn the relationship between the RSV and the outcome of interest in the observational sample, and to “transport” it to the experimental sample.
Assumption 2.2 (Stability of the remotely sensed variable).
@@
    Common support: for some outcome support $\mathcal{Y}$, $\Pr(Y\in\mathcal{Y}\mid S=e,X)=1$ almost surely, and $\Pr(Y=y\mid S=o,X)$ is bounded away from zero almost surely for all $y\in\mathcal{Y}$.
@@
    Coverage: $\Pr(R=r\mid S,X,D)$ is bounded away from zero almost surely for all $r\in\mathcal{R}$.
@@
Assumption 2.2(i) is the main assumption of our framework: the conditional distribution of the remotely sensed variable $R$, given $(X,D,Y)$, is stable across the experimental and observational samples. This allows us to “transport” the measurement error distribution from the observational sample to the experimental sample. Importantly, this condition does not require stability on the underlying treatment effects, which may differ across samples. See Remark 3.1 for comparisons between Assumption 2.2(i) in the RSV model and alternative assumptions in alternative models.
Returning to our two leading examples, Assumption 2.2(i) requires that the conditional distribution of tree cover pixels $R$, given environmental outcomes $Y$ and interventions $D$ (as well as other pre-treatment covariates), is stable across the experimental and the observational samples. Analogously, it requires that the conditional distribution of the satellite image $R$, given village-level poverty $Y$ and the anti-poverty program $D$ (as well as other pre-treatment covariates), is stable across the experimental and observational samples.
@@
Figure 2: Our main assumption (Assumption 2.2(i)) is plausible in real data. We compare $\Pr(R\mid S=e,D=0,Y=0)$ with $\Pr(R\mid S=o,D=0,Y=0)$ in Figure 2(b), and $\Pr(R\mid S=e,D=0,Y=1)$ with $\Pr(R\mid S=o,D=0,Y=1)$ in Figure 2(d), using data from Muralidharan et al., (2023) that we analyze in Section 5. Because the satellite image $R\in\mathbb{R}^{4000}$ is high dimensional, we visualize the density of its standardized first principal component on the right hand side, for units highlighted on the left hand side.
Our main assumption is empirically plausible, as illustrated by Figure 2. We use data from an anti-poverty program in India, where outcomes are observed.
If Assumption 2.2 holds, then the conditional densities of the RSV given the outcome and treatment should be the same across the experimental and observational samples. They appear to coincide in this empirical setting. 777 When $X=\varnothing$, Assumption 2.2 imposes four equalities: $\Pr(R\mid S=e,D=d,Y=y)=\Pr(R\mid S=o,D=d,Y=y)$ for $d\in\{0,1\}$ and $y\in\{0,1\}$. Each equality can be evaluated with a diagnostic plot if outcome data are available. For example, using the experimental and observational units satisfying $D=0$ and $Y=0$ in Figure 2(a), we can visualize whether the density of $R\mid S=e,D=0,Y=0$ aligns with the density of $R\mid S=e,D=0,Y=0$ in Figure 2(b). Since $R$ is high dimensional, we simplify the visualization by comparing the densities of its first principal component.
The remaining aspects of Assumption 2.2 are weak regularity conditions. Assumption 2.2(ii) requires that the outcome in the observational sample has a common (or larger) support than the outcome in the experiment. Assumption 2.2(iii) ensures that the RSV distribution does not degenerate for any stratum. Assumption 2.2(iv) requires that we observe some data from both the experimental and observational samples.
Assumptions 2.1 and 2.2 imply identification when the observational sample has complete cases.
When the observational sample has incomplete cases, we require a further assumption. In other words, if the treatment is missing or deterministic in the observational sample, then a further restriction is necessary for point identification.
Assumption 2.3 (Observational completeness).
@@
Assumption 2.3 imposes only one of two conditions.
Assumption 2.3(i) implies that we have access to complete cases, i.e. some observations of $(X,D,Y,R)$ where $D$ has variation. Within the observational sample, the treatment is observed and varies, although it may suffer from unobserved confounding.
Whenever we have complete cases, under Assumption 2.3(i), no further causal assumptions are needed. In particular, the treatment $D$ may have a direct effect on the remotely sensed variable $R$. In Example  2.1, this would allow the environmental program to affect satellite images both indirectly, i.e. via crop burning, and directly, e.g. via visible investments in farm equipment.
If Assumption 2.3(i) is violated, then we have no complete cases, i.e. no observations of $(X,D,Y,R)$ where $D$ is variable. Without joint observations of the outcome and treatment, a further restriction is needed. Assumption 2.3(ii) fills this gap, requiring that the treatment $D$ only affects the remotely sensed variable $R$ via its effect on the outcome $Y$. In Example 2.1, we may be comfortable assuming that the PES contract has no direct effect on the specific infrared band used to measure charred soil in satellite images.
Assumption 2.3(ii) may become more plausible when $Y$ is a vector of outcomes. Several outcomes may approximate all mechanisms though which the treatment affects the RSV. For readability, we focus on scalar outcomes in the main text, and vector outcomes in Appendix E.
@@
$D$$Y$$R$
Figure 3: Causal graph of Assumptions 2.3(i) versus 2.3(ii). Complete cases allow the dotted line.
Figure 3 illustrates our identifying assumptions as a causal graph. The treatment affects the outcome, which in turn affects the RSV. Depending on which version of Assumption 2.3 is imposed, the treatment may also have a direct effect on the RSV, as illustrated by the dotted line. Table 3 summarizes the implications of our main assumptions.
@@ 3 Main result: Identification
To streamline notation, we initially focus on the setting where Assumption 2.3(ii) holds, then return to the setting where Assumption 2.3(i) holds at the end of this section. We will also assume the outcome is binary, i.e. $\mathcal{Y}=\{0,1\}$. Appendices E and F extend our results to (multivalued) discrete and continuous outcomes, respectively.
@@ 3.1 Current practice may have positive or negative bias
In empirical research, it is common for researchers to use RSVs in two steps: (i) researchers train a predictor of the outcome $Y$ from the remotely sensed variable $R$ in the observational sample; (ii) the predictor is applied on the experimental sample, and its predictions are used as surrogate outcomes to estimate treatment effects. While intuitive, this empirical strategy can lead to arbitrary bias in the downstream causal estimate, i.e. for the ATE in the experimental sample.
Suppose there are no pre-treatment covariates for simplicity. The widely used two-step estimation procedure implicitly targets the estimand $\widetilde{\theta}=\widetilde{\mu}(1)-\widetilde{\mu}(0)$, where $\widetilde{\mu}(d):=\mathbb{E}\left\{\mathbb{E}(Y\mid R,S=o)\mid D=d,S=e\right\}$ for $d\in\{0,1\}$. Within this expression, the first step estimates the conditional expectation function $\mathbb{E}(Y\mid R=r,S=o)$. The second step evaluates and averages this function on the treated and untreated subgroups in the experimental sample.
If the RSV fails to predict the outcome, i.e. if $\mathbb{E}(Y|R,S=o)=\mathbb{E}(Y|S=o)$, then the implicit target $\widetilde{\theta}$ is zero regardless of the true treatment effect. Common practice would return a precise estimate of zero, even though the RSV provides no information about treatment effects in this case.
More importantly, even if the RSV does predict the outcome, the implicit target $\widetilde{\theta}$ can incur biases with arbitrary signs for the ATE in the experimental sample, even if units are randomly allocated between the experimental and observational samples.
Proposition 3.1 (Bias of common practice).
@@
  1. 1.
    The bias is $\widetilde{\theta}-\theta=\widetilde{\mu}(1)-\mu(1)=\mu(1)\int\{w(r)-1\}\Pr(R=r|Y=1,S=e)\mathrm{d}r$, where $w(r)=\frac{\Pr\{Y(0)=1\mid S=e\}\Pr(R=r\mid D=1)}{\Pr\{Y(1)=1\mid S=e\}\Pr(R=r\mid D=0)}.$
  2. 2.
    There exists a data-generating process satisfying the above restrictions with $\widetilde{\theta}-\theta>0$, and a different data-generating process with $\widetilde{\theta}-\theta<0$.
Proposition 3.1 derives the bias of current empirical practice, which uses the RSV as a surrogate outcome. Under Assumptions 2.2 and 2.3(ii), the conditional distribution of the RSV $\Pr\left(R\mid Y,S=o\right)$ is stable across the experimental and the observational samples. Existing practice attempts to transport the predictions $\Pr(Y\mid R,S=o)$ into the experimental sample. By Bayes’ rule, this induces a bias due to differences in the marginal distributions of the outcomes and RSVs across the samples.
Remark 3.1 (Comparison to the surrogacy framework).
@@
The surrogacy framework can be viewed as a version of the moment condition model with auxiliary data studied by e.g. Chen et al., (2005), Chen et al., (2005), and Graham et al., (2016). In our terminology, those works require that the conditional distribution of the outcome given the RSV and treatment is stable across samples. By contrast, we require that the conditional distribution of the RSV given the outcome and treatment is stable across samples. Figure 2 suggests that our assumption is plausible in a real development application with satellite images.
The surrogacy model, and related models which lead to the estimand $\widetilde{\theta}$, have been extended to include negative controls e.g. Ghassami et al., (2022); Imbens et al., (2024). As generalizations of the surrogacy model, they suffer from the same drawback formalized in Proposition 3.1.
Similar to the RSV framework, the single negative control framework (Park et al.,, 2024) has one auxiliary variable. The frameworks have two key differences. Unlike the single negative control model, we allow the treatment $D$ to affect both the outcome $Y$ and the auxiliary variable $R$ when Assumption 2.3(i) holds. When no complete cases exists, i.e. the setting covered by Assumption 2.3(ii), the single negative control model provides no guidance of how to proceed.
@@ 3.2 Example: Current practice may underestimate environmental impacts
We revisit an experiment conducted by Jack et al., (2025) that studied whether payments for ecosystems services (PES) contracts can incentivize farmers to reduce crop burning. We measure the average treatment effect $\theta$ of being offered a PES contract $D\in\{0,1\}$ on the likelihood that a farmer burns their fields $Y\in\{0,1\}$. Here, $Y=1$ means not burning, so a positive $\theta$ means environmental benefit.
It is costly to measure whether the crop residue on a particular field has been burned, requiring a professional surveyor. Therefore, it is natural to turn to a remotely sensed variable $R$ for the outcome $Y$. To construct such a remotely sensed variable, Jack et al., (2025) link surveyor-collected measurements of crop burning to satellite-based spectral indices and then train a supervised learning algorithm to predict whether these fields have been burned. Here, $R\in\{0,1\}$ is a classifier for whether a field has not been burned, which applies a threshold rule to a machine learning prediction of the probability that a field has not been burned. [^8] Jack et al., (2025) construct two binary RSVs for crop burning by applying two alternative threshold rules to the estimated probability a field has not been burned. We use their “max accuracy” RSV in the main text, and we report analogous results in Table 4 using the authors’ “balanced accuracy” RSV.
Our causal assumptions are plausible in this setting. Because burning crops would alter satellite images, but altering satellite images would not burn crops, $R$ is a post-outcome variable rather than a pre-outcome variable. Since access to PES contracts was randomized at the village level, Assumption 2.1 is satisfied by design. Since the authors conducted randomized spot checks, surveying the outcome $Y$ for randomly selected fields, Assumption 2.2 is also satisfied by design, if we define the observational sample $S=o$ as the fields in which these random spot checks occurred. Finally, since specific infrared bands are used to form $R$, it is plausible that PES contracts only affect these specific infrared bands via crop burning, so Assumption 2.3(ii) is reasonable. Figure 10 illustrates the two samples used in our re-analysis.
@@
Notes: The RSV is the field-level “maximum accuracy” label defined by Jack et al., (2025), which applies a threshold rule to the predicted probability of not being burned. The “observational sample” has fields that received a random spot check, and the “experimental sample” has other fields. For illustration, we conduct linear estimation, controlling for stratum fixed effects. Standard errors are based on $5000$ bootstrap replications clustered at the village level.
@@
We implement the common empirical practice: we use the RSV as a surrogate for crop burning. Column (1) of Table 2 estimates $\widetilde{\theta}$. Offering any PES contract to farmers appears to reduce crop burning by 7.9%. [^9] We modify the main specification of Jack et al., (2025) in two ways. First, while they distinguish between two types of PES contracts, we define the treatment as whether any PES contract was offered. Second, the authors’ analyze the effects of the PES contracts by defining a farmer-level outcome, whereas we analyze effects at the field level.
@@
However, by Proposition 3.1, $\widetilde{\theta}$ is typically biased for $\theta$. To quantify the magnitude of its bias, we estimate $\beta:=\mathbb{E}(R\mid Y=1)-\mathbb{E}(R\mid Y=0)$ using information collected through random spot checks of the fields. Under Assumption 2.3(ii), an algebraic argument in Appendix B.1 yields $\widetilde{\theta}=\beta\theta$ in linear settings. Column (2) of Table 2 reports the estimate of $\beta$. It suggests that $\widetilde{\theta}$ understates the treatment effect of being offered any PES contract by approximately 47%.
3.3 Main result: A novel formula

### v3

@@ 2.2 Main Assumption: Stability
[…]
Assumption 1 (Experimental unconfoundedness).
@@
    SUTVA: Y=DY(1)+(1D)Y(0)Y=DY(1)+(1-D)Y(0) almost surely.
@@
In many empirical applications involving RSVs, such as Examples 1 and 2, Assumption 1 is satisfied by design: experimental units are chosen as aggregates without spillovers, e.g., villages, and the treatment is randomly assigned for these experimental units. We defer to Appendix J the study of quasi-experimental settings, such as instrumental variables and difference-in-differences.
Under Assumption 1, if we were to observe the outcome in the experimental sample, the ATE could be identified using standard arguments. However, the outcome is not observed in the experiment; instead, we have an RSV.
We resolve this measurement issue by leveraging the observational sample. Intuitively, the idea is to learn the relationship between the RSV and the outcome of interest in the observational sample and to “transport” it to the experimental sample.
Assumption 2 (Stability of the remotely sensed variable).
@@
    Common support: for some outcome support $\mathcal{Y}$, $\Pr(Y\in\mathcal{Y}\mid S=e,X)=1$ almost surely, and $f_{Y}(y\mid S=o,X)$ is bounded away from zero almost surely for all $y\in\mathcal{Y}$.
@@
    Coverage: $f_{R}(r\mid S,X,D)$ is bounded away from zero almost surely for all $r\in\mathcal{R}$.
@@
Assumption 2(i) is the main assumption of our framework: the conditional distribution of the remotely sensed variable $R$, given $(X,D,Y)$, is stable across the experimental and observational samples. This allows us to “transport” the measurement error distribution from the observational sample to the experimental sample. Importantly, this condition does not require stability of the underlying treatment effects, which may differ across samples. See Remark 2 for comparisons between Assumption 2(i) in the RSV model and assumptions in alternative models.
Returning to our two leading examples, Assumption 2(i) requires that the conditional distribution of tree cover pixels $R$, given environmental outcomes $Y$ and interventions $D$ (as well as other pre-treatment covariates), is stable across the experimental and observational samples. Analogously, it requires that the conditional distribution of the satellite image $R$, given village-level poverty $Y$ and the anti-poverty program $D$ (as well as other pre-treatment covariates), is stable across the experimental and observational samples.
@@
Figure 2: Our main assumption (Assumption 2(i)) is plausible in real data. We compare $f_{R}(R\mid S=e,D=0,Y=0)$ with $f_{R}(R\mid S=o,D=0,Y=0)$ in Figure 2(b), and $f_{R}(R\mid S=e,D=0,Y=1)$ with $f_{R}(R\mid S=o,D=0,Y=1)$ in Figure 2(d), using data from the Smartcard experiment conducted by Muralidharan et al. (2016) that we analyze in Section 5. Because the satellite image $R\in\mathbb{R}^{4000}$ is high-dimensional, we visualize the density of its standardized first principal component on the right hand side, for units highlighted on the left hand side.
Our main assumption is empirically plausible, as illustrated by Figure 2. We use data from an anti-poverty program in India, where outcomes are observed. If Assumption 2 holds, then the conditional densities of the RSV given the outcome and treatment should be the same across the experimental and observational samples. They appear to coincide in this empirical setting. [^8] When $X=\varnothing$, Assumption 2 imposes four equalities: $f_{R}(R\mid S=e,D=d,Y=y)=f_{R}(R\mid S=o,D=d,Y=y)$ for $d\in\{0,1\}$ and $y\in\{0,1\}$. Each equality can be evaluated with a diagnostic plot if outcome data are available. For example, using the experimental and observational units satisfying $D=0$ and $Y=0$ in Figure 2(a), we can visualize whether the density of $R\mid S=e,D=0,Y=0$ aligns with the density of $R\mid S=o,D=0,Y=0$ in Figure 2(b). Since $R$ is high-dimensional, we simplify the visualization by comparing the densities of its first principal component.
The remaining aspects of Assumption 2 are weak regularity conditions. Assumption 2(ii) requires that the outcome in the observational sample has a common (or larger) support than the outcome in the experiment. Assumption 2(iii) ensures that the RSV distribution does not degenerate for any stratum. Assumption 2(iv) requires that we observe some data from both the experimental and observational samples.
Assumptions 1 and 2 imply identification when the observational sample has complete cases. When the observational sample has incomplete cases, we require a further assumption. In other words, if the treatment is missing or deterministic in the observational sample, then a further restriction is necessary for point identification.
Assumption 3 (Observational completeness).
@@
Assumption 3 imposes only one of two conditions.
Assumption 3(i) implies that we have access to complete cases, i.e., some observations of $(X,D,Y,R)$ where $D$ has variation. Within the observational sample, the treatment is observed and varies, although it does not need to be randomized and may suffer from unobserved confounding. Whenever we have complete cases, under Assumption 3(i), no further causal assumptions are needed. In particular, the treatment $D$ may have a direct effect on the remotely sensed variable $R$. In Example  1, this would allow the environmental program to affect satellite images both indirectly, i.e., via crop burning, and directly, e.g., via visible investments in farm equipment.
If Assumption 3(i) is violated, then we have no complete cases, i.e., no observations of $(X,D,Y,R)$ where $D$ is variable. Without joint observations of the outcome and treatment, a further restriction is needed. Assumption 3(ii) fills this gap, requiring that the treatment $D$ affects the remotely sensed variable $R$ only via its effect on the outcome $Y$. In Example 1, we may be comfortable assuming that the PES contract has no direct effect on the specific infrared band used to measure charred soil in satellite images. Assumption 3(ii) may also become more plausible when $Y$ is a vector of outcomes. Several outcomes may approximate all mechanisms through which the treatment affects the RSV. For exposition, we focus on scalar outcomes in the main text and we generalize to vector outcomes in Appendix E.
@@
SSDDYYRR
Figure 3: Causal graph for remotely sensed variables under Assumptions 3(i) versus 3(ii). Complete cases allow the dotted line. Assumption 2 rules out the line from $S$ to $R$.
Figure 3 illustrates our identifying assumptions as a causal graph. The treatment affects the outcome, which in turn affects the RSV. By Assumption 2, the sample indicator does not change the conditional distribution of the RSV given the treatment and outcome. Depending on which version of Assumption 3 is imposed, the treatment may have a direct effect on the RSV, as illustrated by the dotted line. Table A.1 summarizes the implications of our main assumptions.
@@ 3 Main Result: Identification
To streamline notation, we initially focus on the setting where Assumption 3(ii) holds, then return to the setting where Assumption 3(i) holds at the end of this section. For exposition, we will also assume that the outcome is binary, i.e., $\mathcal{Y}=\{0,1\}$. Finally, we assume that the support of $R$ is at least as large as the support of $Y$ (it may be binary, discrete, or continuous). Appendices E and F extend our results to discrete and continuous outcomes, respectively.
@@ 3.1 Current Practice may have Positive or Negative Bias
In empirical research, it is common to use RSVs in two steps: (i) researchers train a predictor of the outcome $Y$ from the remotely sensed variable $R$ in the observational sample, and (ii) the predictor is applied to the experimental sample, where its predictions are used as surrogate outcomes to estimate treatment effects. While intuitive, this empirical strategy can lead to arbitrary bias for the ATE in the experimental sample.
Suppose there are no pre-treatment covariates for simplicity. The widely used two-step estimation procedure implicitly targets the estimand $\widetilde{\theta}=\widetilde{\mu}(1)-\widetilde{\mu}(0)$, where $\widetilde{\mu}(d):=\mathbb{E}\left\{\mathbb{E}(Y\mid R,S=o)\mid D=d,S=e\right\}$ for $d\in\{0,1\}$. Within this expression, the first step estimates the conditional expectation function $\mathbb{E}(Y\mid R=r,S=o)$. The second step evaluates and averages this function over the treated and untreated subgroups in the experimental sample.
As a starting point, if the RSV fails to predict the outcome, i.e., if $\mathbb{E}(Y|R,S=o)=\mathbb{E}(Y|S=o)$, then the implicit target $\widetilde{\theta}$ is zero regardless of the true treatment effect. Common practice would return a precise estimate of zero, even though the RSV provides no information about treatment effects in this case.
More importantly, even if the RSV does predict the outcome, the implicit target $\widetilde{\theta}$ can incur bias with arbitrary sign for the ATE in the experimental sample.
Proposition 1 (Bias of common practice).
@@
  1. i.
    The bias is $\widetilde{\theta}-\theta=\mu(1)\int\{w(r)-1\}f_{R}(r|Y=1,S=e)\mathrm{d}r$, where $w(r)=\frac{\Pr\{Y(0)=1\mid S=e\}f_{R}(r\mid D=1)}{\Pr\{Y(1)=1\mid S=e\}f_{R}(r\mid D=0)}.$
  2. ii.
    There exists a data-generating process satisfying the above restrictions with $\widetilde{\theta}-\theta>0$, and a different data-generating process satisfying the above restrictions with $\widetilde{\theta}-\theta<0$.
Proposition 1 derives the bias of current empirical practice, which uses the RSV as a surrogate outcome. Under Assumptions 2 and 3(ii), the conditional distribution of the RSV $f_{R}\left(R\mid Y,S=o\right)$ is stable across the experimental and observational samples. Existing practice attempts to transport the predictions $f_{Y}(Y\mid R,S=o)$ into the experimental sample. By Bayes’ rule, this induces a bias due to differences in the marginal distributions of the outcomes and RSVs across the samples. Importantly, the bias of common practice arises even if units are randomly allocated into the experimental and observational samples. [^9] See Appendix B for a formal verification of this claim.
Remark 1 (Bias summary).
To gain insight into the source of bias, we consider three nested cases. (i) In the extreme “irrelevance” case, the RSV fails to predict the outcome. We would like an estimator whose variance is infinity. However, $\mathbb{E}(Y|R,S=o)$ is a constant function of $R$, so $\widetilde{\theta}=0$; the common practice always reports zero. (ii) In the simple linear case, something similar occurs. Suppose the data-generating process satisfies $\mathbb{E}(Y|R,S=o)=\tilde{\beta}_{0}+\tilde{\beta}R$ and $\mathbb{E}(R|Y,D,S=e)=\beta_{0}+\beta Y$. Combining these expressions, $\widetilde{\theta}=\tilde{\beta}\beta\theta_{0}$; the common practice suffers from attenuation bias (see Appendix B for further details). (iii) In the general case, where the relationships among variables are nonlinear, Proposition 1 shows that the bias of common practice can be positive or negative.
Remark 2 (Comparison to the surrogacy framework).
@@
The surrogacy framework can be viewed as a version of the moment condition model with auxiliary data studied by e.g., Chen et al. (2005, 2008); Graham et al. (2016). In our terminology, those works require that the conditional distribution of the outcome given the RSV and treatment is stable across samples. By contrast, we require that the conditional distribution of the RSV given the outcome and treatment is stable across samples. Figure 2 suggests that our assumption is plausible in a real development application with satellite images.
The surrogacy model, and related models which lead to the estimand $\widetilde{\theta}$, have been extended to include negative controls (e.g., Ghassami et al., 2022; Imbens et al., 2024). As generalizations of the surrogacy model, they suffer from the same drawback formalized in Proposition 1.
Similar to the RSV framework, the single negative control framework (Park et al., 2024) has one auxiliary variable. The frameworks have two key differences. Unlike the single negative control model, we allow the treatment $D$ to affect both the outcome $Y$ and the auxiliary variable $R$ when Assumption 3(i) holds. When no complete cases exist, i.e. the setting covered by Assumption 3(ii), the single negative control model provides no guidance of how to proceed.
@@ 3.2 Example: Current Practice may Underestimate Environmental Impacts
We revisit an experiment conducted by Jack et al. (2025) that studied whether payments for ecosystem services (PES) contracts can incentivize farmers to reduce crop burning. We measure the average treatment effect $\theta$ of being offered a PES contract $D\in\{0,1\}$ on the likelihood that a farmer burns their fields $Y\in\{0,1\}$. Here, $Y=1$ means not burning, so a positive $\theta$ means environmental benefit.
It is costly to measure whether the crop residue on a particular field has been burned, requiring a surveyor to make frequent visits to rural fields. Therefore, it is natural to turn to a remotely sensed variable $R$ for the outcome $Y$. To construct such a remotely sensed variable, Jack et al. (2025) link surveyor-collected measurements of crop burning to satellite-based spectral indices and then train a supervised learning algorithm to predict whether these fields have been burned. Here, $R\in\{0,1\}$ is a classifier for whether a field has not been burned, which applies a threshold rule to a machine learning prediction of the probability that a field has not been burned. [^10] Jack et al. (2025) construct two binary RSVs for crop burning by applying two alternative threshold rules to the estimated probability a field has not been burned. We use their “max accuracy” RSV in the main text, and we report analogous results in Table G.1 using the authors’ “balanced accuracy” RSV.
Our causal assumptions are plausible in this setting. Because burning crops would alter satellite images, but altering satellite images would not burn crops, $R$ is a post-outcome variable rather than a pre-outcome variable. Since access to PES contracts was randomized at the village level, Assumption 1 is satisfied by design. Since the authors conducted randomized spot checks, surveying the outcome $Y$ for randomly selected fields, Assumption 2 is also satisfied by design if we define the observational sample $S=o$ as the fields in which these random spot checks occurred. Finally, since specific infrared bands are used to form $R$, it is plausible that PES contracts affect these specific infrared bands only via crop burning, so Assumption 3(ii) is reasonable. Figure G.1 illustrates the two samples used in our re-analysis.
@@
Notes: The RSV is the field-level “maximum accuracy” label defined by Jack et al. (2025), which applies a threshold rule to the predicted probability of not being burned. The “observational sample” has fields that received a random spot check, and the “experimental sample” has other fields. For illustration, we conduct linear estimation, controlling for stratum fixed effects. Standard errors are based on $5000$ bootstrap replications clustered at the village level.
@@
We implement the common empirical practice: we use the RSV as a surrogate for crop burning. Column (1) of Table 2 estimates $\widetilde{\theta}$. Offering any PES contract to farmers appears to reduce crop burning by 7.9%. [^11] We modify the specification of Jack et al. (2025) in two ways. (i) While they distinguish between two types of PES contracts, we define the treatment as whether any PES contract was offered. (ii) They analyze the effects of PES contracts by defining a farmer-level outcome, whereas we analyze effects at the field level.
@@
However, by Proposition 1, $\widetilde{\theta}$ is typically biased for $\theta$. To quantify the magnitude of its bias, we estimate $\beta:=\mathbb{E}(R\mid Y=1)-\mathbb{E}(R\mid Y=0)$ using information collected through random spot checks of the fields. Under Assumption 3(ii), an algebraic argument in Appendix B yields $\widetilde{\theta}=\beta\theta$ in this setting. [^12] Compared to Remark 1(ii), here we have $\mathbb{E}(Y|R,S=o)=R$, which implies $\tilde{\beta}=1$ and hence $\widetilde{\theta}=\beta\theta_{0}$. Column (2) of Table 2 reports the estimate of $\beta$. It suggests that $\widetilde{\theta}$ understates the treatment effect of being offered any PES contract by approximately 47%. This motivates us to ask: what should empirical researchers do instead?
3.3 Main Result: An Identification Formula